What does 3i^33-2i^27 equal and how do you find the answer?
Answer by stanbon(75887)

What does 3i^33-2i^27 equal and how do you find the answer?
-----------
Notice:
i^1 = i
i^2 = -1
i^3 = -i
i^4 = 1
----------
then the cycle repeats itself over and over again
----------------
So i^33 = [i^32]i = [1]^8 i = i
And i^27= [i^24]i^3 = 1^6*(-i) = -i
---------------
Your problem:
3*i^33 -2i^27
= 3i -2*(-i)
= 3i+2i
= 5i
==================
Cheers,
Stan H.